The perimeter of a rectangular yard is 66m and its area is 270 square meters. Find the dimensions of the yard.
Answer by checkley77(12569)

xy=270
x=270/y
2x+2y=66
2(270/y)+2y=66
540/y+2y=66
(540+2y^2)/y=66
540+2y^2=66y
2y^2-66y+540=0
2(y^2-33y+270)=0
2(y-15)(y-18)=0
y-15=0
y=15 x=18 ans.
y-18=0
y=18 x=15 ans.
Proofs:
15*18=270
270=270
2*15+2*18=66
30+36=66
66=66